### v1

# Dark energy as a geometrical effect in geodetic brane gravity

Abstract
Within the framework of the modified geodetic brane gravity, conformed by the Regge-Teitelboim model and enhanced with a linear term in the extrinsic curvature of the brane, the possibility that under an FRW geometry this theory emulates the so-called dark energy is discussed. The cosmological behavior of this model displays a self-(non-self)-accelerated expansion of this universe which is caused by a combination of usual matter and gravitational geometric effects controlled by a $\beta$ parameter that accompanies the correction $K$ term. Indeed, the self-accelerated branch, provided by the trace $K$ model raises the question of whether the extrinsic curvature correction terms might be suitable for dark energy candidates. We discuss the analytical expression obtained for $\rho_{\text{\tiny dark}}$ in addition to the main cosmological parameters such as the state parameter $\omega_{\text{\tiny eff}}$ and the deceleration parameter $q$. Moreover, when we call for the contribution of dark radiation-like energy to be switched off, $\Omega_{dr,0}\to 0$, we find the same acceleration behavior, as well as the same dark energy content provided by the DGP theory.
I Introduction
Any attempt to solve the still unexplained mystery of the dark matter/energy, as long as it is based on fundamental geometric symmetries which play a central role in respected gravitational theories, is welcome. The dark energy nature constitutes a tantalizing question to which the cosmology and the particle physics are attempting to provide a compelling answer. Following this line of reasoning, within the geodetic brane gravity based in the Regge-Teitelboim (RT) model, we would like to explore a possible explanation to such a puzzle where extrinsic curvature terms can provide findings that reproduce many of the characteristics that exhibits the so-called dark matter/energy. The RT gravity theory generalizes General Relativity (GR) in a particular theoretical direction [1]. The theory is based on a simple particle/string-inspired assumption that our Universe is a 4-dimensional surface isometrically embedded, and evolving, in a higher-dimensional flat ambient spacetime. In this view, our 4-dimensional spacetime is regarded as the trajectory (worldvolume) of a 3-dimensional extended object, probing a higher-dimensional flat ambient space. Within this assumption, one must not overlook the local isometric embedding theorem [2, 3]. The RT action is identical to the Einstein-Hilbert (EH) action of GR but, with the crucial difference that the embedding functions $X^{\mu}(x^{a})$ rather than the metric $g_{ab}$ defined of the worldvolume, are elevated to the level of field variables. The field equations result in the Einstein ones but contracted with the extrinsic curvature tensor of the embedding which represents a generalization of the Einstein equations in the sense that the RT equations are fulfilled for every solution of the Einstein equations. The theory has attractive properties such as the already mentioned built-in Einstein limit and in particular, within the braneworld context, a dark companion to any primitive energy density. Under the names of geodetic brane gravity or embedding gravity, the RT model has been widely investigated by many authors as an interesting formulation to describe gravity, cosmological implications in extra dimensions as well as a suitable alternative for quantization of gravity [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
The trace $K$ model, a linear term in the extrinsic curvature, represents an alternative term such as RT gravity based on the isometrical embeddings that might underlie cosmic acceleration, and which does not alter the order of the equation of motion [19]. In this spirit, the dynamics of the brane is still driven by the Gauss-Codazzi, Codazzi-Mainardi, and Ricci integrability conditions of the geometry of surfaces [20, 21]. Certainly, the trace $K$ term offers some singular features that deserve full interest. For instance, within the braneworld cosmological scenario, this model acts like a fluid, emulating effects such as that of a cosmological constant, and a self-(non-self)-acceleration dynamical behavior at late times and producing unnatural matter.
In this paper, within the extended objects framework, we attempt to go one step beyond pure RT gravity and reignite the discussion regarding the dark energy companion of the primitive $\rho$ for the RT cosmology as discussed in [22, 23, 24], by exercising the geometrical features contained in the trace $K$ model. Specifically, in this regard, we show that a kind of dark energy is a geometrical effect. Certainly, some braneworld models at low energies have become interesting scenarios providing late-time acceleration behaviors without the need to consider dark energy by adding in the extra mysterious matter. We observe that $\beta$, the parameter accompanying the trace $K$ model, is the only input parameter in our development that produces this effect. In this sense, with the claim of exploring the source of the dark energy, within the modified gravity theories context, one of the requirements is the built-in Einstein limit. On the basis of geodetic brane gravity, it is interesting and motivating to improve the theory by incorporating properly pathologies-free geometric terms so that physical effects from other more elaborated theories can be reproduced. Despite the bulk being flat Minkowski, there are local effects on the brane via conserved quantities associated with the Poincaré invariance, and the correction $K$ term which can be thought of as a remnant of the existence of a 5-dimensional EH action. We find that our proposal, when applied to a homogeneous and isotropic universe that has an FRW geometry, leads to a suggestive modification of the Friedmann equation which includes an additional term, $\rho_{\text{\tiny dark}}$, in the effective energy density and that it is treated in the manner that the primitive $\rho$. The presence of such $\rho_{\text{\tiny dark}}$ originates mainly from the extrinsic curvature correction term $K$ and the matter confined to the braneworld. The trace $K$ model requires careful observation; needles to say such a term resembles the Gibbons-Hawking-York (GHY) counterterm but with the essential difference that in our approach $K$ contains time derivatives of the field variables that contribute to the dynamics of the brane. It is worth mentioning this model has cropped up in other contexts ranging from lipid membranes to particle physics and brane-world gravity [25, 26, 27].
In the first part of our work, our strategy is, under an FRW geometry on the brane, to make use of the invariance under reparametrizations of the model and identify conserved quantities that help to integrate the equation of motion, and thus lead us to identify a Friedmann type equation. By invoking, and imposing, the traditional point of view of the cosmology for the Friedmann equation, we may identify an effective energy density where $\rho_{\text{\tiny dark}}$ is a companion piece of the primitive energy density $\rho$. Indeed, we can express the dark energy contribution in terms of the associated extrinsic curvature correction parameter, $\beta$, and the confined matter, producing a self-(non-self)-acceleration behavior of this type of universe. One could wonder if the answer to the tantalizing question of dark energy is related to the inherent extrinsic curvature and its embedding properties, as was perceived in [28], so our approach suggests that is the answer. Subsequently, we analytically compute important cosmological parameters such as the equation of state parameter and the deceleration parameter. These confirm the existence of a type of dark energy density and the existence of a self-(non-self)-acceleration behavior through the emergence of Big Chill and Big Bounce histories of this universe. It is worth mentioning that this model suffers from a fine-tuning problem, so one must rely on numerical analysis.
This paper is organized as follows. In Sec. II we review the geometrical aspects of the geodetic brane gravity enhanced with the trace $K$ model. We discuss the resulting equation of motion, the built-in Einstein limit, and the definition of the concept of embedding matter. Additionally, we accommodate our development to an FRW geometry. In Sec. III we obtained an analytical expression for the dark energy contribution in our framework as well as some cosmological parameters of interest. Further, we specialize our analysis to the DGP approach when the contribution from the extra dimension is turned off. Finally, in Sec. IV we conclude and discuss our findings.
II Modified Regge-Teitelboim gravity
The modified Regge-Teitelboim gravity, [19], including a matter content, is provided by
$S[X^{\mu}]=\int_{m}d^{3+1}x\,\sqrt{-g}\,\left(\frac{\alpha}{2}\mathcal{R}+\beta\,K+L_{\text{\tiny matt}}\right),$ (1)
where $\alpha$ and $\beta$ are constants with appropriate dimensions. The scalars $\mathcal{R}$ and $K$ are formed using the induced metric $g_{ab}$ and the extrinsic curvature $K_{ab}$, that encode the geometrically most important derivatives of the embedding functions $X^{\mu}$,
$g_{ab}$ $=$ $\eta_{\mu\nu}\partial_{a}X^{\mu}\partial_{b}X^{\nu},$ (2)
$K_{ab}$ $=$ $-\eta_{\mu\nu}n^{\mu}\nabla_{a}\partial_{b}X^{\nu}.$ (3)
Here, $\partial_{a}X^{\mu}=:e^{\mu}{}_{a}$ and $n^{\mu}$ represent the tangent vectors and the normal vector to the world volume $m$, respectively, and $\nabla_{a}$ is the covariant derivative compatible with $g_{ab}$, and $\eta_{\mu\nu}$ is the flat Minkowski metric in the background spacetime. $\mu=0,1,\ldots,4$ and $a=0,1,2,3$ (see Appendix A for details about of the geometrical constructions behind the description of extended objects). Certainly, $\mathcal{R}$ is the Ricci scalar defined on $m$ (our Universe) obtained from $g_{ab}$, and $K=g^{ab}K_{ab}$ denotes the trace of $K_{ab}$ where $g^{ab}$ stands for the inverse of $g_{ab}$. Further, $g=\det(g_{ab})$. The possible matter fields coupled to $m$ are embodied in the Lagrangian $L_{\text{\tiny matt}}$.
The action (1) resembles Einstein-Hilbert action defined on $m$, augmented with the dynamical term $K$, but with the difference that the independent variables are the embedding functions $X^{\mu}(x^{a})$, where $x^{a}$ are local coordinates of $m$. In this sense, the RT gravity is, by construction, a continuation of string theory so it justified that the field variables are the $X^{\mu}$ and not the induced metric components $g_{ab}$.
The variational principle applied to $S$, under infinitesimal changes $X^{\mu}(x^{a})\rightarrow X^{\mu}(x^{a})+\delta X^{\mu}(x^{a})$, leads to the fact that the physical deformations of the action are the normal to $m$, $\delta_{\perp}S=0$[21, 29, 30] so that the only field equation in a coordinate-invariant form is
$\left(\alpha G^{ab}+\beta T_{\text{\tiny K}}^{ab}-T^{ab}_{\text{\tiny m}}\right)K_{ab}=0,$ (4)
where $G_{ab}=\mathcal{R}_{ab}-\frac{1}{2}g_{ab}\mathcal{R}$ is the worldvolume Einstein tensor, $T_{ab}^{\text{\tiny K}}:=K_{ab}-g_{ab}K$, and $T_{ab}^{\text{\tiny m}}:=\frac{(-2)}{\sqrt{-g}}\frac{\partial(\sqrt{-g}\,L_{\text{\tiny matt}})}{\partial g^{ab}}$ is the worldvolume energy-momentum tensor. Although the action (1) is of second order in derivatives of $X^{\mu}$, equation (4) represents a second-order in derivatives equation of motion. In passing, by using the contracted Codazzi-Mainardi equation, (40), one verify that $\nabla_{a}T_{\text{\tiny K}}^{ab}=0$, in analogy with the usual Bianchi identity $\nabla_{a}G^{ab}=0$.
With support on the invariance under reparametrizations of $m$, that is the symmetry of the action (1), the mechanical content of $m$ can also be obtained from the current conservation law, $\nabla_{a}f^{a\,\mu}=0$, [29], with
$f^{a\,\mu}=-\left(\alpha G^{ab}+\beta T_{\text{\tiny K}}^{ab}-T^{ab}_{\text{\tiny m}}\right)e^{\mu}{}_{b},$ (5)
being the conserved stress tensor which is purely tangential to the world volume. Indeed, $f^{a\,\mu}$ obeys
-- $\nabla_{a}\left(\alpha G^{ab}+\beta T_{\text{\tiny K}}^{ab}-T^{ab}_{\text{\tiny m}}\right)e^{\mu}{}_{b}$
$+$ $\left(\alpha G^{ab}+\beta T_{\text{\tiny K}}^{ab}-T^{ab}_{\text{\tiny m}}\right)K_{ab}\,n^{\mu}=0.$
From this, we immediately infer the equation of motion (4) as well as the fact that $T^{ab}_{\text{\tiny m}}$ is covariantly conserved, $\nabla_{a}T^{ab}_{\text{\tiny m}}=0$.
Some comments are in order. The scalar $K$ that appears in the action (1) is similar to the GHY counterterm in pure GR but this analogy needs a word of caution. Within the framework of extended objects, floating in an ambient spacetime, the $K$ term possesses time derivatives of the variables $X^{\mu}$, see (3); hence, such a term provides dynamics on the hypersurface $m$. Regarding this, $T_{\text{\tiny K}}^{ab}$ is a purely geometric tensor that plays the role of a energy-momentum-like tensor. In this line of reasoning the $K$ term can be thought of as a matter type contribution to the Einstein-Hilbert term as we will see below. On the other hand, in view of (4), we introduce
$\mathsf{T}^{ab}:=\alpha G^{ab}+\beta T_{\text{\tiny K}}^{ab}-T^{ab}_{\text{\tiny m}}.$ (6)
This entails that the equation of motion (4) can be written in the form
$\left[G^{ab}-\frac{1}{\alpha}\left(T^{ab}_{\text{\tiny m}}-\beta T_{\text{\tiny K}}^{ab}\right)\right]K_{ab}=0,$ (7)
or,
$\nabla_{a}\left(\mathsf{T}^{ab}\,\partial_{b}X^{\mu}\right)=0.$ (8)
Therefore, these imply that, on the one hand, if $\mathsf{T}^{ab}=0$ with $\beta=0$, every solution of Einstein equations is automatically a solution of the Regge-Teitelboim gravity; on the other hand, if $\mathsf{T}^{ab}\neq 0$, different solutions are allowed to this type of gravity. Regarding this last point, as discussed in [23, 31], $\mathsf{T}^{ab}$ must be understood as an additional fictional matter contrasting with the ordinary one, and labelled as dark matter or embedding matter, making it a possible explanation for dark matter. In connection with this, notice that
$\nabla_{a}(T^{ab}_{K}\,\partial_{b}X^{\mu})=0,\qquad\text{or}\qquad\partial_{a}\left(\sqrt{-g}\,j^{\mu\,a}_{K}\right)=0,$ (9)
with the geometric current $j^{\mu\,a}_{K}$ defined by
$j^{\mu\,a}_{K}:=\left(K^{ab}-K\,g^{ab}\right)\partial_{b}X^{\mu}.$ (10)
II.1 Friedmann cosmology in modified RT gravity
We will assume that our universe is the evolution of a $3$-dimensional extended object, say $\Sigma$, in a $5$-dimensional Minkowski spacetime parametrized in spherical coordinates, $ds^{2}_{5}=-dt^{2}+da^{2}+a^{2}d\Omega^{2}_{3}$, where $d\Omega^{2}_{3}=d\chi^{2}+\sin^{2}\chi d\theta^{2}+\sin^{2}\chi\sin^{2}\theta d\phi^{2}$ denotes the unit $3$-sphere. At this stage a parametric representation for this universe $m$, assuming homogeneity, isotropy and a closed topology is provided by
$x^{\mu}=X^{\mu}(x^{a})=(t(\tau),a(\tau),\chi,\theta,\phi),$ (11)
where $x^{\mu}$ are the local coordinates for the ambient spacetime and $\tau$ denotes the proper time for an observer at rest concerning $m$. Recall, $a,b=0,1,2,3$. The orthonormal basis adapted to $m$ is provided by the four tangent vectors $e^{\mu}{}_{a}=\partial X^{\mu}/\partial x^{a}$ complemented with the unit spacelike normal vector $n^{\mu}=(1/N)(\dot{a},\dot{t},0,0,0))$ where $N:=\sqrt{\dot{t}^{2}-\dot{a}^{2}}$. Here and henceforth, an overdot denotes a $\tau$-derivative. The metric induced on $m$ from the ambient spacetime, (2), is
$ds^{2}_{4}=g_{ab}dx^{a}dx^{b}=-N^{2}d\tau^{2}+a^{2}d\Omega^{2}_{3},$ (12)
while the non-trivial components of the extrinsic curvature (3) are
$K^{\tau}{}_{\tau}=\frac{\dot{t}^{2}}{N^{3}}\frac{d}{d\tau}\left(\frac{\dot{a}}{\dot{t}}\right),\quad\quad K^{\chi}{}_{\chi}=K^{\theta}{}_{\theta}=K^{\phi}{}_{\phi}=\frac{\dot{t}}{\dot{Na}}.$ (13)
The Gauss-Codazzi equations (39), allows us to know the world volume Riemann and Ricci tensors, and the Ricci scalar, in terms of the extrinsic curvatures. Hence,
$K$ $=$ $\frac{1}{N^{3}}(\dot{t}\ddot{a}-\dot{a}\ddot{t})+\frac{3\dot{t}}{Na},$ (14)
$\mathcal{R}$ $=$ $\frac{6\dot{t}}{N^{4}a}(\dot{t}\ddot{a}-\dot{a}\ddot{t})+\frac{6\dot{t}^{2}}{N^{2}a^{2}}.$ (15)
It is required that $T^{ab}_{\text{\tiny m}}$ acquire the form of a general perfect fluid
$T^{ab}_{\text{\tiny m}}=\left(\rho+P\right)\eta^{a}\eta^{b}+P\,g^{ab},$ (16)
with energy density $\rho(a)$ and pressure $P(a)$, and $\eta^{a}$ being the timelike unit normal vector to the extended object at a fixed $\tau$.
The non-vanishing components of the conserved tensors $T^{ab}_{\text{\tiny m}}$, $T_{\text{\tiny K}}^{ab}$ and $G_{ab}$ can be computed for the Friedmann-like geometry in a straightforward manner
$T_{\text{\tiny m}}^{\tau}{}_{\tau}$ $=-\rho$ $\qquad T_{\text{\tiny m}}^{\chi}{}_{\chi}$ $=T_{\text{\tiny m}}^{\theta}{}_{\theta}=T_{\text{\tiny m}}^{\phi}{}_{\phi}=P,$ (17)
$T_{\text{\tiny K}}^{\tau}{}_{\tau}$ $=-\frac{3\dot{t}}{Na}$ $\qquad T_{\text{\tiny K}}^{\chi}{}_{\chi}$ $=T_{\text{\tiny K}}^{\theta}{}_{\theta}=T_{\text{\tiny K}}^{\phi}{}_{\phi}=-\frac{\dot{t}^{2}}{N^{3}}\frac{d}{d\tau}\left(\frac{\dot{a}}{\dot{t}}\right)-\frac{2\dot{t}}{Na},$
$G^{\tau}{}_{\tau}$ $=-\frac{3\dot{t}^{2}}{N^{2}a^{2}}$ $\qquad G^{\chi}{}_{\chi}$ $=G^{\theta}{}_{\theta}=G^{\phi}{}_{\phi}=-\frac{2\dot{t}^{3}}{N^{4}a}\frac{d}{d\tau}\left(\frac{\dot{a}}{\dot{t}}\right)-\frac{\dot{t}^{2}}{N^{2}a^{2}}.$
As discussed in [19], for example, the flat and hyperbolic geometries can be included by convenient generalizations of the embedding (11). To accommodate these geometries in the cosmological development for this model, we just need to do the promotion $\dot{t}\longrightarrow(\dot{a}^{2}+k)^{1/2}$ with $k=-1,0,1$, accompanied by the usual cosmic gauge with a function $N\longrightarrow\mathcal{N}$, specific to each embedding.
By suitably writing the equation of motion (18), we wish to address the analysis of the mentioned embedding matter equation of motion (8) when $a=\tau$ and $\mu=t$ and obtain
$\frac{1}{\sqrt{-g}}\partial_{\tau}\left(\sqrt{-g}\,\mathsf{T}^{\tau\tau}\,\partial_{\tau}X^{t}\right)=0.$ (18)
This is accompanied by the energy-momentum conservation law $\nabla_{a}T^{ab}_{\text{m}}=0$ which produces the integrability condition $\dot{\rho}+3\left(\frac{\dot{a}}{a}\right)(\rho+P)=0$. In view of this, from the geometric conservation law $\nabla_{a}T^{ab}_{K}=0$, one could expect a cosmological energy-momentum conservation equation related to the trace $K$ model. We wish to address the solving analysis of these equations to explore the dark matter content provided by this brane model, as we will see shortly.
III Dark energy from modified geodetic brane cosmology
The solution provided by (18) leads to a constant of motion. To obtain this we first insert the values (17) into the definition (6) to get $\mathsf{T}^{\tau\tau}=\frac{3\alpha\dot{t}^{2}}{N^{4}a^{2}}+\frac{3\beta\dot{t}}{N^{3}a}-\frac{\rho(a)}{N^{2}}$. Similarly, by considering $g=-N^{2}a^{6}\sin^{4}\chi\sin^{2}\theta$ and then inserting these results into expression (18) we get
$\partial_{\tau}\left[\frac{3\alpha\,a\,\dot{t}^{3}}{N^{3}}+\frac{3\beta\,a^{2}\dot{t}^{2}}{N^{2}}-\frac{a^{3}\,\rho(a)\,\dot{t}}{N}\right]=0.$ (19)
After a trivial integration followed by choosing the cosmic gauge, $N=1$, and the inclusion of the three main geometries by $\dot{t}\longrightarrow(\dot{a}^{2}+k)^{1/2}$, we arrive at our master equation
$-3\alpha a\,(\dot{a}^{2}+k)^{3/2}-3\beta a^{2}\,(\dot{a}^{2}+k)+\rho a^{3}\,(\dot{a}^{2}+k)^{1/2}=:\frac{\omega}{\sqrt{3\,\alpha}},$ (20)
with $\omega$ being a constant, the fingerprint of the background spacetime. In fact, $\omega$ plays a double duty. On the one hand, it is related to the conserved bulk energy and, on the other hand, this parameterizes the deviation from the Einstein limit when $\omega=0$ in addition to $\beta=0$. In fact, after some algebra and solving an involved cubic equation, (20) leads to an effective Friedmann equation [19].
If we do not renounce the point of view of the standard cosmology, ignoring the existence of an extra dimension and the brane gravity, we would rearrange (20) so that it can be cast in the form
$\dot{a}^{2}+k=\frac{1}{3\alpha}\left(\rho+\rho_{\text{\tiny dark}}\right)a^{2},$ (21)
where $\rho_{\text{\tiny dark}}$ encode all the extra contributions, physical and geometrical, provided by the extra dimension to the primitive energy density $\rho$. In this sense, the evolution of this type of universe is dictated by an effective energy density $\rho_{\text{\tiny eff}}:=\rho+\rho_{\text{\tiny dark}}$ and not solely by the primitive energy density $\rho$. Further, the Friedmann-like equation (21) can be written in the fashion $\dot{a}^{2}+k=\frac{1}{3\alpha}f(\rho)a^{2}$ once we determine the form of $\rho_{\text{\tiny dark}}$ as we will see shortly. It is worth mentioning that the form (21) can also be obtained by assuming $T^{ab}_{K}$ has a contribution similar to that of a perfect fluid of the form $T_{K}^{\tau}{}_{\tau}=\rho_{K}$ and $T_{K}^{i}{}_{i}=P_{K}$, with $i=\chi,\theta,\phi$ along with the requirement $\dot{\rho}_{K}+3\frac{\dot{a}}{a}(\rho_{K}+P_{K})=0$, [16].
If (20) and (21) are in accordance, $\rho_{\text{\tiny dark}}$ results in the root of a cubic equation. Certainly, plugging (21) into (20) yields
$3\beta_{*}\,(\rho+\rho_{\text{\tiny dark}})+\frac{\omega}{a^{4}}=-\rho_{\text{\tiny dark}}\left(\rho+\rho_{\text{\tiny dark}}\right)^{1/2},$ (22)
where we have put $\beta_{*}:=\frac{\beta}{\sqrt{3\,\alpha}}$. By defining $\mathcal{Z}:=(\rho+\rho_{\text{\tiny dark}})^{1/2}$ and substituting it into the latter equation, followed by a reorganization, we find
$\mathcal{Z}^{3}+3\,\beta_{*}\,\mathcal{Z}^{2}-\rho\,\mathcal{Z}+\frac{\omega}{a^{4}}=0.$ (23)
This is precisely the implicit equation for the dark matter content that we intend to analyze. In passing, we see that for $\beta_{*}=0$ we get $\rho_{\text{\tiny dark}}^{2}(\rho+\rho_{\text{\tiny dark}})=\frac{\omega^{2}}{a^{8}}$ in complete agreement with the result found in [22, 23, 24]. By invoking the change of variable $\varrho:=\mathcal{Z}+\beta_{*}$ in (23) we find the depressed cubic equation $\varrho^{3}+p(a)\,\varrho+q(a)=0$ where
$p=-\left(\rho+3\,\beta_{*}^{2}\right),\qquad\text{and}\qquad q=\beta_{*}\left(\rho+2\,\beta_{*}^{2}\right)+\frac{\omega}{a^{4}}.$ (24)
The solution for the depressed cubic equation can be written in terms of trigonometric or hyperbolic functions with the use of a couple handy identities, $4\cos^{3}\theta-3\cos\theta=\cos(3\theta)$ or $4\cosh^{3}\theta-3\cosh\theta=\cosh(3\theta)$. Certainly, depending on the values chosen for functions (24) we could have solutions with physical interest. In any case, the general forms of the solutions are provided by
$\rho_{\text{\tiny dark}}=\left[2\sqrt{\frac{|p|}{3}}\cosh\left(\frac{1}{3}\cosh^{-1}(\phi)\right)-\beta_{*}\right]^{2}-\rho,$ (25)
or
$\rho_{\text{\tiny dark}}=\left[2\sqrt{\frac{|p|}{3}}\cos\left(\frac{1}{3}\cos^{-1}(\phi)+n\frac{2\pi}{3}\right)-\beta_{*}\right]^{2}-\rho,$ (26)
with $n=0,1,2$. Here, $\phi:=-\frac{q}{2}\left(-\frac{3}{p}\right)^{3/2}$. More precisely, (25), applies when $p\leq 0$ and $|\phi|\geq 1$, while (26), will be the one of interest when $p\leq 0$ and $|\phi|\leq 1$. Additionally, if $p>0$ we would have another possible scenario; in such a case the main function will be $\sinh$ instead of $\cosh$ or $\cos$. We confine ourselves to the physical solutions provided by (25) or (26). Note that whatever the physical solution is, the definite positivity of the total energy density holds, which is one of the exclusive features of the geodetic brane gravity [24]. Notice in passing that this conclusion holds even for $\rho_{\text{\tiny dark}}<0$. Additionally, since $T^{ab}_{K}$ is covariantly conserved, and insisting on the cosmological principle, we assume that $P_{\text{\tiny dark}}$ can be obtained from (25) or (26)
$P_{\text{\tiny dark}}=-\left(\frac{a}{3\dot{a}}\,\dot{\rho}_{\text{\tiny dark}}+\rho_{\text{\tiny dark}}\right).$ (27)
We find convenient to rewrite expressions (25) and (26) in terms of energy density parameters. Indeed, for simplicity, from now on we restrict ourselves to consider dust matter content on the brane so that $\rho(a):=\frac{\rho_{0}}{a^{3}}$, with $\rho_{0}$ being a constant. When introducing the Hubble parameter $H:=\frac{\dot{a}}{a}$ as well the energy density parameters
$\Omega_{m,0}:=\frac{\rho_{0}}{3\alpha H_{0}^{2}},\quad\Omega_{\beta,0}:=\frac{\beta}{\alpha\,H_{0}},\quad\Omega_{dr,0}:=\frac{\omega}{(3\alpha)^{3/2}H_{0}^{3}},$ (28)
where as usual, $H_{0}$ is the Hubble constant, (25) and (26) turn into the expressions
$\frac{\rho_{\text{\tiny dark}}}{\alpha H_{0}^{2}}=-\frac{3\Omega_{m,0}}{a^{3}}+\frac{1}{3}\left\{2\sqrt{\Omega_{\beta,0}^{2}+\frac{3\Omega_{m,0}}{a^{3}}}\cosh\left[\frac{1}{3}\cosh^{-1}\left(-\frac{\Omega_{\beta,0}\left(\Omega_{\beta,0}^{2}+\frac{9}{2}\frac{\Omega_{m,0}}{a^{3}}\right)+\frac{27}{2}\frac{\Omega_{dr,0}}{a^{4}}}{\left(\Omega_{\beta,0}^{2}+\frac{3\Omega_{m,0}}{a^{3}}\right)^{3/2}}\right)\right]-\Omega_{\beta,0}\right\}^{2},$ (29)
and
$\frac{\rho_{\text{\tiny dark}}}{\alpha H_{0}^{2}}=-\frac{3\Omega_{m,0}}{a^{3}}+\frac{1}{3}\left\{2\sqrt{\Omega_{\beta,0}^{2}+\frac{3\Omega_{m,0}}{a^{3}}}\cos\left[\frac{1}{3}\cos^{-1}\left(-\frac{\Omega_{\beta,0}\left(\Omega_{\beta,0}^{2}+\frac{9}{2}\frac{\Omega_{m,0}}{a^{3}}\right)+\frac{27}{2}\frac{\Omega_{dr,0}}{a^{4}}}{\left(\Omega_{\beta,0}^{2}+\frac{3\Omega_{m,0}}{a^{3}}\right)^{3/2}}\right)+n\frac{2\pi}{3}\right]-\Omega_{\beta,0}\right\}^{2},$ (30)
where $n=0,1,2$. The choice of one of the solutions is in dependence of the values of the density energy parameters (28) satisfying the inequalities provided by the discriminant of the cubic equation. Indeed, for the depressed cubic equation, the discriminant is given by $\Delta=-\left(4p^{3}+27q^{2}\right)$. If $\Delta<0$ we shall have one real root provided by (29) and two non-real complex conjugate roots while that for $\Delta>0$ we shall have three distinct real roots provided by (30). Considering (24) and (28), it is straightforward to show that for
$\frac{\Omega_{\beta,0}\left(\Omega_{\beta,0}^{2}+\frac{9}{2}\frac{\Omega_{m,0}}{a^{3}}\right)+\frac{27}{2}\frac{\Omega_{dr,0}}{a^{4}}}{\left(\Omega_{\beta,0}^{2}+\frac{3\Omega_{m,0}}{a^{3}}\right)^{3/2}}>1,\qquad\quad\text{or}\qquad\quad\frac{\Omega_{\beta,0}\left(\Omega_{\beta,0}^{2}+\frac{9}{2}\frac{\Omega_{m,0}}{a^{3}}\right)+\frac{27}{2}\frac{\Omega_{dr,0}}{a^{4}}}{\left(\Omega_{\beta,0}^{2}+\frac{3\Omega_{m,0}}{a^{3}}\right)^{3/2}}<1,$ (31)
we have one physical root or three physical ones, respectively.
In Figure (1) we depict $\rho_{\text{\tiny dark}}$ with $n=0$, for illustration, for some chosen values of the $\Omega$s. In both cases, we observe that, while $\rho(a)$ explodes as $a\longrightarrow 0$, the effective energy density remains finite and positive. Notice also that, for some specific choice of the energy density values, $\rho_{\text{\tiny dark}}$ decreases slower than the primitive energy $\rho\thicksim 1/a^{3}$.
Figure 1: The effective $\rho_{\text{\tiny dark}}$ contained in the possible brane closed universe with some parameters choices. The last plot is a comparison between $\rho_{\text{\tiny dark}}$ and the primitive energy $\rho\thicksim 1/a^{3}$
We remark in passing that, the non-standard Friedmann equation in our approach is provided by $\dot{a}^{2}+k=f(a;\Omega_{k,0},\Omega_{m,0},\Omega_{\beta,0},\Omega_{dr,0})a^{2}$ where the energy density parameters are given in (28), and we have additionally introduced the energy density parameter $\Omega_{k,0}:=-\frac{k}{H_{0}^{2}}$[19]. One can go one step further and cast this Friedmann-like equation in the suggestive form $\dot{a}^{2}+U(a;\Omega_{s})=0$ where $U$ stands for an effective potential parameterized by the density energies $\Omega_{s}$, (28), [19]. This fashion is convenient since it allows the system to be faced as an equivalent one-dimensional non-relativistic mechanical problem in which the total mechanical energy is zero. Certainly, we have
$\frac{U}{H_{0}^{2}}=-\Omega_{k,0}-\frac{a^{2}}{9}\left\{2\left(\Omega_{\beta,0}^{2}+\frac{3\Omega_{m,0}}{a^{3}}\right)^{1/2}F\left[\frac{1}{3}F^{-1}\left(-\frac{\Omega_{\beta,0}\left(\Omega_{\beta,0}^{2}+\frac{9}{2}\frac{\Omega_{m,0}}{a^{3}}\right)+\frac{27}{2}\frac{\Omega_{dr,0}}{a^{4}}}{\left(\Omega_{\beta,0}^{2}+\frac{3\Omega_{m,0}}{a^{3}}\right)^{3/2}}\right)\right]-\Omega_{\beta,0}\right\}^{2},$ (32)
where $F=\cosh x$ for $|x|>1$, or $F=\cos x$ for $|x|\leq 1$. In passing, regarding the values of the energy densities (28), we emphasize that these are restricted to obey the normalization condition
$\left(1-\Omega_{k,0}\right)^{1/2}\left(\Omega_{k,0}+\Omega_{m,0}-1\right)-\Omega_{\beta,0}\left(1-\Omega_{k,0}\right)=\Omega_{dr,0}.$ (33)
This formula will prove very useful to carry out numerical analysis as well as analyze special limits. Indeed, for $\Omega_{dr,0}=\Omega_{\beta,0}=0$ together we immediately recover the usual FRW cosmology whereas for only $\Omega_{dr,0}=0$, we recover the DGP approach admitting the two well-known branches of the theory, [32, 33, 19]. We depict this potential for some chosen values of the energy densities in Figure (2). In both cases we observe
Figure 2: The effective potential $U$ describing possible brane trajectories for a closed universe and some parameters choices.
two general features: the singularity of $U(a,\Omega_{s})$ at small scales, $a\to 0$, as well as at large scales $a\to\infty$. More precisely, in dependence on the chosen parameters we find either a universe expanding forever (Big Chill), which is a signal of the presence of mysterious dark energy in order to counteract the gravitational interaction or else the universe may present a Big Bounce destiny, that is, in a state of constant expansion (Big Bang) and eventually reverses and re-collapses (Big Crunch). In Figure (3) we have a graph enclosing different possibilities for the universe expansion in our approach depending upon the densities $\Omega_{\beta,0}$ and $\Omega_{dr,0}$.
Figure 3: We depict several possibilities for the expansion of this closed universe in dependence of $\Omega_{\beta,0}$ and $\Omega_{m,0}$. The bluish-purple color region is associated with a Big Crunch evolution of the universe, and the coral color region corresponds to a Big Bounce evolution whereas the blue area is in accordance with a universe with a Big Chill behavior. The yellow region corresponds to non-physical universes.
Having disclosed the unconventional behavior of the dark energy content in our approach, in accordance with our symmetry considerations we can assume the resulting effective fluid in (21) to have an equation of state of the form $P_{\text{\tiny eff}}=\omega_{\text{\tiny eff}}\rho_{\text{\tiny eff}}$, where $\omega_{\text{\tiny eff}}$ is subject to the values of the $\Omega$s. Indeed, by assuming $P_{\text{\tiny eff}}=-\left[\frac{a}{3\dot{a}}\left(\dot{\rho}+\dot{\rho_{\text{\tiny dark}}}\right)+\left(\rho+\rho_{\text{\tiny dark}}\right)\right]/\left(\rho+\rho_{\text{\tiny dark}}\right)$, using (27), (30) with $n=0$ for simplicity, as well as the redshift $z$ expression when the universe was $1/(1+z)$ of its present size, it follows that
$\omega_{\text{\tiny eff}}=-1+\frac{3(1+z)^{3}}{\omega_{m,\beta,z}^{2}\Phi_{m,\beta,dr,z}}$$\left\{\Omega_{m,0}(\Omega_{\beta,0}+\Phi_{m,\beta,dr,z})\right.$ (34)
$\left.+\frac{3(1+z)\omega_{m,\beta,z}\left[(1+z)^{2}\Omega_{m,0}^{2}\Omega_{\beta,0}+\Omega_{dr,0}[(1+z)^{3}\Omega_{m,0}-\frac{8}{3}\Omega_{\beta,0}^{2}]\right]\sin(\varphi_{m,\beta,dr,z})}{\sqrt{\omega_{m,\beta,z}^{6}-\Omega_{m,\beta,dr,z}^{2}}}\right\},$
where
$\Omega_{m,\beta,dr,z}$ $:=-\Omega_{\beta,0}\left[\Omega_{\beta,0}^{2}+\frac{9}{2}(1+z)^{3}\Omega_{m,0}\right]$ (35)
$-\frac{27}{2}(1+z)^{4}\Omega_{dr,0},$
$\omega_{m,\beta,z}$ $:=\left[\Omega_{\beta,0}^{2}+3(1+z)^{3}\Omega_{m,0}\right]^{1/2},$
$\varphi_{m,\beta,dr,z}$ $:=\frac{1}{3}\cos^{-1}\left(\frac{\Omega_{m,\beta,dr,z}}{\omega_{m,\beta,z}}\right),$
$\Phi_{m,\beta,dr,z}$ $:=-\Omega_{\beta,0}+2\omega_{m,\beta,z}\cos\left(\varphi_{m,\beta,dr,z}\right).$
It is worthwhile to note that our results specialize to the GR limit for $\Omega_{dr,0}=\Omega_{\beta,0}=0$. Indeed, for such a case we find $\Omega_{m,0,0,z}=0,\omega_{m,0,z}=\sqrt{3(1+z)^{2}}\,\Omega_{m,0},\varphi_{m,0,0,z}=\pi/6$ and $\Phi_{m,0,0,z}=\sqrt{3}\,\omega_{m,0,z}$, so that, with these values we obtain $\omega_{\text{\tiny eff}}=0$ as in the standard cosmology. One can show that for $\beta=0$, at least numerically, the state parameter $\left.\omega_{\text{\tiny eff}}\right|_{\beta=0}>-\frac{1}{3}$ for allowed physical regions, that is, we do not have acceleration behavior. In some sense, this explains the reason why the non-vanishing values of $\Omega_{dr,0}$ and $\Omega_{\beta,0}$ induce self-(non-self)-acceleration for this type of universe at late times. In Figure (4) we plot the state parameter for some chosen values of the energy densities $\Omega_{s}$. In comparison with the standard cosmology, we observe the presence of dark energy content because $\omega_{\text{\tiny eff}}<0$.
Figure 4: The state parameter is depicted for some values. The graph on the left corresponds to a Big Chill behavior, describing a self-acceleration universe. On the contrary, the right plot corresponds to a Big Bounce dynamics.
We further observe that for the chosen values, and at the present time, we do not have a complete energy content to fulfill the condition $-1<\omega_{\text{\tiny eff}}<-1/3$. It is mandatory to choose either another set of energy density values or the inclusion of some matter fields, in order to get a self-acceleration behavior to reproduce the known observations for the dark energy. Additionally, it is worthwhile to observe that at late times the combination of the energy density values mimics, in some sense, the role of a cosmological constant. This is so because we are able the reach the value $\omega_{\text{\tiny eff}}=-1$ in our plot, which is a characteristic of a de Sitter type universe. Further, the state parameter never crosses the effective phantom line, $\omega_{\text{\tiny eff}}$. As a consequence, with the chosen values $\Omega_{s}$, the usual dominant energy condition is not violated in agreement with the non-negative property of $\rho_{\text{\tiny eff}}$.
Another parameter worth exploring is the deceleration parameter, $q(a)=-1-\dot{H}/H$. After a lengthy but straightforward computation, we find that
$q=-1+\frac{9(1+z)^{2}}{\Phi_{m,\beta,dr,z}^{2}+9(1+z)^{2}\Omega_{k,0}}$$\left\{\Omega_{k,0}+\frac{(1+z)}{2}\frac{\Phi_{m,\beta,dr,z}}{\omega_{m,\beta,z}^{2}}\left[\Omega_{m,0}(\Omega_{\beta,0}+\Phi_{m,\beta,dr,z})\right.\right.$ (36)
$-\left.\left.\frac{2\omega_{m,\beta,z}\left[\left(4(1+z)\Omega_{dr,0}+\Omega_{m,0}\Omega_{\beta,0}\right)\omega_{m,\beta,z}^{2}+\Omega_{m,0}\Omega_{m,\beta,dr,z}\right]\sin(\varphi_{m,\beta,dr,z})}{\sqrt{\omega_{m,\beta,z}^{6}-\Omega_{m,\beta,dr,z}^{2}}}\right]\right\}$
where we have considered the structures (35). We observe immediately that the transition value when the deceleration turned into acceleration is provided by $\Phi_{m,\beta,dr,z}=0$. Additionally, the GR limit is reached for $\Omega_{dr,0}=\Omega_{\beta,0}=0$, that is, $q=\frac{\Omega_{m,0}/a^{3}}{2(\Omega_{m,0}/a^{3}+\Omega_{k,0}/a^{2})}$, as usual. In figure (5) we depict the $q$ deceleration parameter for some chosen values of the $\Omega_{s}$. There, for a Big Chill expansion we found a typical $q$ deceleration behavior where the transition redshift $z_{0}$, for some given values for the $\Omega_{s}$, is $z_{0}\approx-0.419$. On the contrary, for a Big Bounce expansion we observe that the interval $-0.52<z<0.95$ does not allow a physical scenario which is in agreement with the interval $21.2142.09<a<$ 21.214 found for the corresponding potential plot in Figure (2).
Figure 5: The deceleration parameter is depicted for the values $\Omega_{\beta,0}=-0.5$, $\Omega_{dr,0}=-0.1$, $\Omega_{m,0}=2.5$ and $\Omega_{k,0}=-0.761$. For the left plot, the transition value is given when $z\approx-0.419$, which corresponds to the value $a\approx 1.72$ (See figure (2) for comparison.).
III.1 DGP-like dark energy from modified RT cosmology
The radiation-like energy density, $\Omega_{dr,0}$, represents the fingerprint of the extra-dimensional ambient spacetime. As we already discussed, it influences the dynamics of this brane-like universe. At the limit $\Omega_{dr,0}=0$, the master equation (20) reduces remarkably to $03\alpha(\dot{a}^{2}+k)+3\beta\,a(\dot{a}^{2}+k)^{1/2}-\rho\,a^{2}=$ 0. When inserting (21) and recalling that $\rho_{\text{\tiny eff}}=\rho+\rho_{\text{\tiny dark}}$, we have that $\rho_{\text{\tiny eff}}=\rho-\beta_{*}\sqrt{\rho+\rho_{\text{\tiny dark}}}$. We determine immediately that
$\rho_{\text{\tiny dark}}=-3\beta\sqrt{H^{2}+\frac{k}{a^{2}}}.$ (37)
This special limit assures that the effective dark energy within this geodetic brane context is nothing more than a geometric effect provided by the correction term $K$. Obviously, we can include a cosmological constant in our development but we claim to highlight the role of the inclusion of the trace $K$ model. Additionally, from this expression, we immediately identify that $\beta$ parameter plays the role of the inverse of the crossover scale of the DGP approach [32, 34]. Nonetheless, a word of caution is needed; although $\beta$ can take positive and negative values, it is mandatory that $\rho_{\text{\tiny dark}}$ be subject to the condition $\rho_{\text{\tiny eff}}\geq 0$. In Figure (6) we depict $\rho_{\text{\tiny dark}}$ for this special case.
Figure 6: The dark energy $\rho_{\text{\tiny dark}}$ for the DGP approach considering some values of the energy densities. In both cases, we compare the behavior of $\rho_{\text{\tiny dark}}$ with the primitive energy density $\rho\sim 1/a^{3}$.
IV Concluding remarks
In this paper, we have analyzed the dark energy content within the geodetic brane gravity context enhanced with an extrinsic curvature term associated with the shape of the brane. We provide evidence that such dark energy can originate from the correction of extrinsic curvature terms appearing in the action functional that models our Universe. To show this we have considered a FRW geometry defined onto our braneworld probing a $5$-dimensional Minkowski background spacetime. Our approach is free of a cosmological constant with the idea of highlighting the role that the correction term $K$ plays in development. In this sense, it plays the role of a cosmological constant providing self-(and non-self) acceleration behavior. One can think of the $K$ term as a relic or a possible fingerprint of the existence of a $5$-dimensional EH action that disappears at some time. By assuming a traditional point of view for the Friedmann equation, we observed that the regular matter is accompanied by an atypical piece, termed $\rho_{\text{\tiny dark}}$. In our development, the intriguing dark energy effect results as a geometrical achievement originated by the model under consideration. What is notable is to observe how the correction terms associated with the extrinsic curvature of the brane can reproduce many of the general characteristics of the so-called dark energy.
Our approach is not intended to be realistic, but on the contrary, it aims to be the promoter of extrinsic curvature corrections that help to understand several physical phenomena ranging from condensed soft matter to cosmology and particle physics. Because of this, perhaps the inclusion of some matter fields, such as the cosmological constant, can take on a more realistic meaning. A consequent extension of our work presented here is the incorporation of other attractive extrinsic curvature terms such as those provided by the so-called Lovelock brane gravity theory. For instance, the Gibbons-Hawking-York-Mayers type term, $K^{3}-2KK_{ab}K^{ab}+3K^{a}{}_{b}K^{b}{}_{c}K^{c}{}_{a}$, [30]. We believe that such an addition could improve or enrich the cosmological findings in this braneworld context. This will be reported elsewhere.
Acknowledgements.
The authors are grateful to R. Cordero, J. A. Méndez Zavaleta and M. Cruz for discussions and helpful suggestions. ER acknowledges encouragement from ProDeP-México, CA-UV-320: Álgebra, Geometría y Gravitación. JCNZ acknowledges support from a CONAHCYT-México graduate fellowship. GC acknowledges support from a Postdoctoral Fellowship by Estancias Posdoctorales por México 2023(1)-CONAHCYT . Also, ER thanks partial support from Sistema Nacional de Investigadores, México.
Appendix A Extended objects geometry
Consider a $(p+1)$-dimensional manifold $m$ that represents the worldvolume, $m$, of a spacelike brane $\Sigma$. $m$ is embedded in a $N$-dimensional flat Minkowski spacetime, $\mathcal{M}$, with metric $\eta_{\mu\nu}=\text{diag}(-1,1,...,1)$ ($\mu,\nu=0,1,2,\ldots,N-1$). The worldvolume is described by the embedding functions $X^{\mu}(x^{a})$ where $x^{a}$ are local coordinates for $m$ ($a,b=0,1,2,\ldots p$). The tangent vectors to $m$ are given by $e^{\mu}{}_{a}:=\partial X^{\mu}/\partial x^{a}$ which allow us to define an induced metric on $m$ as $g_{ab}:=\eta_{\mu\nu}e^{\mu}{}_{a}e^{\nu}{}_{b}$. As usual, $g^{ab}$ denotes the inverse of $g_{ab}$, and $g=\det(g_{ab})$. Additionally, $m$ is assumed to be time-like so $g<0$. The $i$th normal vector to $m$, $n^{\mu\,i}$, ($i,j=1,2,...,N-p-1$), is defined by the relations $n^{i}\cdot e_{a}=0$ and $n_{i}\cdot n_{j}=\delta_{ij}$. These expressions define the normal vectors up to a sign and a local $O(N-p-1)$ rotation. This gauge freedom requires a gauge field, known as twist potential, given by $\omega_{a}{}^{ij}=-n^{i}\cdot\partial_{a}n^{j}$[21]. In the case of a hypersurface embedding, $i=1$ and the extrinsic twist vanishes identically. Further, $\nabla_{a}$ denotes the (torsionless) covariant derivative compatible with $g_{ab}$.
For codimension one case, ($i=1$), the extrinsic curvature of $m$ is $K_{ab}=-n\cdot\nabla_{a}\nabla_{b}X$, and the mean extrinsic curvature as its trace, $K=g^{ab}K_{ab}$. The corresponding Gauss-Weingarten equations are
$D_{a}e^{\mu}{}_{b}$ $=\Gamma^{c}_{ab}e^{\mu}{}_{c}-K_{ab}n^{\mu},$ (38)
$D_{a}n^{\mu}$ $=K_{ab}g^{bc}e^{\mu}{}_{c},$
where $D_{a}=e^{\mu}{}_{a}D_{\mu}$ with $D_{\mu}$ is the covariant derivative compatible with a possible background metric, say $\mathcal{G}_{\mu\nu}$, and $\Gamma^{a}_{bc}$ are the connection coefficients compatible with $g_{ab}$. The intrinsic and extrinsic geometries of the worldvolume $m$ are related by some integrability conditions. In this sense, the Gauss-Codazzi, and the Codazzi-Mainardi integrability conditions are
00 0 $=$ $\mathcal{R}_{abcd}-K_{ac}K_{bd}+K_{ad}K_{bc},$ (39)
00 0 $=$ $\nabla_{a}K_{bc}-\nabla_{b}K_{ac}.$ (40)